# Unpacking Personal(?!) Health Informatics: An Investigation of Awareness, Understanding, & Leveraged Utility in India

### v1

@@ 3.5. Positionality
The authors of this study acknowledge their positionality as researchers from the Indian cultural context. While some authors still live in India and bring an insider perspective, others live abroad with a background in distributed computing. This diverse positionality allowed for a comprehensive analysis considering local nuances and broader global trends in wearable technology use. The authors also recognize that their interpretations are influenced by their own experiences with using PHI and backgrounds, which may shape the analysis and conclusions drawn from the data. Throughout the research process, efforts were made to remain reflexive and mindful of potential biases, ensuring that the participants’ voices were accurately represented and that the analysis remained grounded in the data.
@@ 4. Findings
Using data triangulation from the various components of the study, we examined the perceptions, usage patterns, motivations, and barriers associated with PHI through an in-depth analysis. The thematic analysis revealed several key themes and sub-themes that offer valuable insights into how users interact with PHI, their trust in the technology, and their challenges/barriers to adoption.
4.1. Collective Health Management Through Sharing Practices
4.1.1. Health Seen as a Collective Responsibility
While major research has focused on individualistic healthcare models, the reflection of Indian practices indicates that health and wellbeing are shared/collective in nature. People often view the health of their family members as a shared responsibility, driving the need to keep on top of each other’s health statuses. For example, in our survey, over 60% (55/87) of participants were already tracking their health using a wearable, and in the interviews, about 45% (10/22) were gifted their tracking device to be used within families to track vital metrics, enabling a collaborative care model.
They viewed health management as a shared, collective process rather than an individual responsibility. They often rely on digital PHI not only to monitor their health but also to care for family members and close contacts. This shared approach fosters mutual support, accountability, and informed decision-making among trusted social circles. For instance, P8 disclosed that apart from getting a personal tracker, he also wanted to track his parents’ vitals to check on them even when living away.
”I also got a fitbit watch for my parents, and my parents wear them. So I can track heart rate and other metrics. So first, I bought for them, and then I got one for myself.” - P8 (M, 26)
Even among people who leave the country for work, their needs based on cultural values lead them to seek ways to track their parents’ health and support their health management remotely. This feeling was expressed through CP12’s personal experience of wanting to monitor his father’s health when he was living abroad a decade ago, but finding a lack of cost-effective means to do so.
”During my initial career days (living abroad), as my dad got surgery, I wanted to be able to monitor him regularly. But there was no cheap and reliable means to do so. So, I had to rely on relatives to inform me of his status. If there was an economical way to monitor my father continuously, it would have helped me a lot.” - CP12 (M, 40)
4.1.2. Understanding the Aspects of Sharing
Through the secondary interviews, co-design sessions on sharing preferences, and a Sharing Awareness Survey, we discovered that individuals are open to sharing their PHI with spouses and doctors without reservation, but are unaware of existing mechanisms to share. They had varied opinions on how they would like their data to be shared and used. Some couples (2/6) in the co-design session were open to their data being used by others as long as they were informed of the use case, while all the rest wanted control over not only sharing but also the storing and usage of the data. The participants’ interest in viewing all kinds of health information about themselves and their family members in a single place showcases a need for a unified system.
The level of openness with sharing decreased in the following order among the participants: Spouses Doctors Parents Siblings Friends Employer Insurance Companies. Some couples (2/6) also mentioned that they wanted to avoid sharing with employers and insurance companies if that was the option.
”So, I am open to sharing everything with my spouse. Maybe just an update on how I am doing with my parents maybe. I don’t think I’ll be sharing much with my siblings or friends except if there are any emergency alerts. I am also open to sharing everything with doctors as long as it makes sense for the doctor to be seeing those for the purpose of the visit.” - CP1 (F, 34)
When inquired about their knowledge and usage of any current sharing frameworks, most of the couples (5/6) were unaware of those already in the market through prominent brands (Apple Health and Samsung Health). This general lack of awareness and usage of such sharing frameworks was also reinforced through the outcome of the survey responses (94/114) from the health sharing feature survey. The aware couple said they were already using it (however sparingly) for things that can be tracked through their wearable. However, they indicated they are unhappy about the lack of control over what can be shared, how it can be shared, and when it can be shared. This observation hints that either the existing platforms are so rigid that users are unhappily using them as little as possible, or they are unaware of it because it is not popular in the urban Indian context. A couple even drew insights from their experience living abroad and the system they have seen being used.
@@ 4.1.2. Understanding the Aspects of Sharing
Another insightful observation among participants interested in sharing was that people with children are interested in tracking and sharing PHI only when there is no overhead of pulling all things together by themselves. While the motivation to track using wearables differs individually (even within a couple), they all were motivated for the overall PHI for themselves and their children if they could do so using a single device.
After incorporating the feedback and inputs, when evaluating the rapid figma prototype, the participants liked the overall functionality of the interface and provided further suggestions. They felt that the prototype mimicked the discussion on their sharing preferences and appreciated the control over them. Some participants also suggested that further information or health tips could be added about the benefits of good sleep and (for women) steps to take when periods are late.
@@
The responses from the Sharing Feature Awareness Survey also indicate a need for a reliable medium of sharing information among Android users and a better demonstration of capabilities among Apple/iPhone users.
”I usually share health reports with family members as and when we get health checkups done, however, they usually get scattered all over in phone storage. If they can be stored all together in one place, that would be a helpful feature. ” - (F, 33)
”I thought sharing was about putting it out on social media platforms. I never knew about sharing features that send health updates to our family or health care provider.” - (M, 24)
Overall, participants’ motivations and needs indicate a need for a single dedicated application or a sharing mechanism where they can pool all kinds of health information about themselves and their trusted circles while maintaining full control over the preferences, such that they do not have additional cognitive load on managing it. While some prefer textual representation of this information, others are comfortable with graphical, while a very limited few are okay with numbers and medical jargon.
4.2. Navigating Health Information in a Fragmented Digital Ecosystem
4.2.1. Unreliable Tools and Systems
Users encounter a patchwork of digital health information, often leaving them uncertain about what to trust. While basic functionalities (like step counting and heart rate monitoring) are widely accepted, there is notable skepticism around advanced metrics. Participants frequently cross-check data with multiple devices, healthcare professionals, and even traditional methods to validate accuracy. This verification process underscores the challenge of navigating a fragmented digital health landscape.
Even the participants who have used various trackers throughout their evolution noted that the current capabilities of PHI systems are not on par with their current purpose or motivation. They emphasized the lack of reliable sources for health information or even the capability of systems to explain their health in understandable terms.
”So one of my first smartwatches was a Pebble smartwatch. It is very old. It started on Kickstarter (—) So that is how I got into it. (—) But more lately. It’s been to serve the purpose of. OK, how do I understand my body a little better? (—) So what it started with and what it is right now, it’s very different. The current systems cannot explain it well to me.” - P19 (M, 36)
4.2.2. Island of Ecosystems
Many of these PHI trackers have their own ecosystems that are specialized for a specific brand or product line, focused on boosting their unique functionalities to promote their line of products. This causes a lack of cross-functional capabilities. This was echoed during the interviews with doctors with different specialties, the common recurring theme was the existing use of specialized trackers for different chronic conditions like diabetes, digital pressure cuffs for hypertension, or even mood journaling for psychiatry. The doctors from endocrinology and medicine confirmed using existing specialized smart devices in practice, albeit their price factor limited it to a few patients. The high financial investment required for reliable devices further compounds these reservations for individuals as well.
Wearables, an integral part of PHI for many users, had technical issues apart from their limitations of cross-functionality on what they could offer the user, causing further frustrations among participants. One such instance is inaccurate heart rate measurements needing recalibration during exercise. This observation shows that mechanisms such as comparison tools or validation indicators that help users assess the accuracy of their health data are currently lacking.
”It happens quite a lot, especially in the gym when I can feel that my heart is like pounding (laughs); like help!. OK, so I can feel my heart beating really fast. But when I see my watch, it will show readings like 75 or 76, which is kind of normal. (—) Then, then I have to again, again like reconfigure that (manually calibrate the reading). Then you put, then you press on that measure. So some kind of calibration is required.” - P3 (F, 33)
Participants noted the lack of reliable health information sources for them to verify their PHI and the reduced level of interoperability among available devices. They expressed that they do not have verified, reputable health information sources to compare their metrics and rely on doctor visits to make sense of them.
Overall, the observations with participants’ struggles to make sense of the data indicate that the fragmented ecosystems are complex and lack inter-compatibility to use them all cohesively for PHI. Simplifying the complexity of digital health data through intuitive design and user-friendly explanations that boost both digital and health literacy could aid in their journey better, as noted by a few participants.
4.3. Digital PHI as a Motivational Tool for Proactive Health Management
Digital PHI catalyzes behavior change by transforming passive data collection into proactive health management. Users appreciate features that provide real-time feedback, set personalized goals, and track progress over time. These tools encourage people to adopt healthier lifestyles, offering actionable insights and reminders. However, the motivational impact is closely tied to individuals’ priorities and preferences. For example, some participants were mindful of their lifestyle, like P16 (F, 26) and P7 (M,32), who track their steps and maintain an active lifestyle during sedentary workdays.
@@ 4.3. Digital PHI as a Motivational Tool for Proactive Health Management
”So I was walking very fast. Then it (the tracker) told me you are walking outdoors; you have walked so much distance. So then, I wish to walk more.” - (P7, M, 32)
Integrating reminders, gamified challenges, and progress-tracking features that motivate users to stay engaged with their health data was among the recommendations from participants in the interviews and rapid prototyping. They considered them resourceful motivators, a means to make healthy habits, or a means for a more personalized health management tool by delivering tailored advice or actionable insights that resonate with their individual and family health goals.
”So the (health) reminder thing, right, it can be a good guilt-tripping system so. (—) more like a motivator (to work out more). This would be a good, good thing to see, like a progress of myself.” - (P17, M, 26)
Participants were eager to view summarized stats over day, week, and month. However, they also wished for them to be presented alternately to increase their engagement. Participants noted that being handy could also be another reason for adopting digital PHI as a reliable tool. This finding indicates the potential of digital PHI as a motivational tool if different presentation modalities are provided.
4.4. Preference for Traditional Methods and Perception of Non-necessity
4.4.1. Traditional and Cultural Influence
Some participants remain cautious about fully adopting digital PHI, favoring traditional health management methods. Concerns about privacy, technological complexity, and physical discomfort (e.g., frequent charging needs and skin irritation) contribute to this resistance. Some users (5) trust their bodily awareness and conventional healthcare practices, finding digital tracking unnecessary or overwhelming.
”I do not prefer to constantly check my miband. I only use it when I run or perform any exercise. I do not find the need to check the history or track, as if something is wrong, I could feel it in my body.” - P21(M,27)
Similarly, P20 (M, 29) emphasized his belief that his body would alert him if something were wrong, reducing the perceived need for a smartwatch and confidence in his healthier lifestyle.
”Whenever you know I’m like whenever I’m sick, (laughs) maybe at the time, something is wrong and I don’t, as you said (tracking health metrics regularly) maybe I’m not much bothered about, the specifics. (—) If there is something wrong with my body, maybe, I know or maybe my body will let me know if I’m feeling something bad. So in that case I will take medication or precaution or something, whatever that is needed at that time.” - P20 (M,29)
Some participants even recognized and reasoned about the disjointedness between the customs of their parents’ generation and the current one, indicating a need for better integrating digital PHI with traditional cultural practices to promote active, healthy living.
”If you think about it, our lifestyles has been completely different from our parents. These (PHI) trackers only work for us. Designers are completely ignoring the population that still leads a more self-reliant lifestyle. If they are able to integrate regular daily activities (like cleaning the house, washing clothes, gardening) as part of measuring other metrics for health markers, I believe it can improve the adoption and create a more proactive health management mindset.” - P15 (F,25)
While participants who seek a more body-over-technology approach exist, they also make recommendations in ways that can improve for a wider audience. This observation shows that the adoption is dictated by the value it offers, which does not meet their needs, and indicates room for improvement.
4.4.2. Barriers or Challenges Causing Digital PHI Resistance
Participants considered an assurance of precision in both basic and advanced features a necessity to maintain trust and encourage ongoing proactive engagement. They compared their childhood experience with traditional measuring devices as a frame of reference for building trust in digital PHI systems, indicating a need to get used to such digital systems.
Comfort and lifestyle fit have also emerged as critical factors influencing the continued use of smart wearables as care devices, causing the aspects of using a wearable to become a barrier for a few. For example, P9 (F, 29), who received a smartwatch as a gift, eventually found the device’s frequent charging needs and complexities overwhelming, not integrating seamlessly into her life.
”So, so one, once I basically invested in this watch. I mean it was all fancy and all of it (—) What I realized is that this was too much of a task for me. The most painful thing is to kind of charge the device. (—) I wanted to do something to kind of keep my health in check, but it got too much for me.” - P9 (F,29)
For others, the physical discomfort or rashes due to wearable material can detract from their overall utility, and others prefer different modalities such as rings, bands, or other wearables to a watch so they can wear their traditional watches, which they enjoy wearing.
”For me, specifically, I do not prefer to wear it continuously because I get rashes from the strap if I wear it for too long regardless of the kind of strap.” - P22 (F,27)
”I’m also trying to go away from a watch format and more a band or a ring or something else of that sort. (—) Mostly because I also wanna like continue wearing my other watches.” - (P19, M, 36)
Overall, the findings indicate that digital PHI can offer options that allow for both proactive tracking and reactive monitoring, acknowledging users’ comfort with conventional health cues would be welcome among users. They also recommend cost-effective solutions to ease the transition from traditional to digital health management.
5. Discussion

### v2

@@ 3.5. Positionality
The authors of this study acknowledge their positionality as researchers from the Indian cultural context. While some authors still reside in India, bringing an insider’s perspective, others live abroad with families in India. This diverse positionality enabled a comprehensive analysis that considered both local nuances and broader global trends in the use of PHI for proactive care. The authors also recognize that their interpretations are influenced by their own experiences with using PHI and backgrounds, which may shape the analysis and conclusions drawn from the data. Throughout the research process, efforts were made to remain reflexive and mindful of potential biases, ensuring that the participants’ voices were accurately represented and that the analysis remained grounded in the data.
@@ 4. Findings
We examined the perceptions, usage patterns, motivations, and barriers associated with through an in-depth analysis. The thematic analysis revealed several key themes and sub-themes that offer valuable insights into their trust in PHI for proactive care, their barriers/enablers to proactivity, and how users interact with PHI for proactive care.
4.1. Barriers to Proactivity
Through the study’s various components, participants identified five recurring barriers at different levels of an individual’s ecosystem that prevent PHI from enabling earlier action and restricting its adoption or prolonged use. Each barrier reduces either the perceived usefulness of continuous sensing or the user’s ability to interpret and act on signals.
4.1.1. Ecosystem fragmentation and interoperability
At the system level, the silo of ecosystems made the process of choosing the tools for PHI arduous. Participants described PHI as scattered across multiple vendor apps and formats. Clinical documents and device data are not consolidated into a single longitudinal view. This fragmentation increases cognitive load and erodes the perceived value of monitoring, since users cannot easily aggregate trends or produce easy-to-read summaries. Several participants explicitly requested unified timelines and export templates (e.g., EHR compatibility) that would enable the actionable use of PHI in clinical encounters.
Another byproduct of fragmentation is also the burden of purchasing wearables being an investment. As each ecosystems have its own compatibilities, any individual user needs to invest effort, time, and money to figure out which ones to adopt to make full use of proactive action. This is evident from P3’s experience, as she explains the process of buying a smartwatch as akin to a risky investment. If a user decides to purchase a higher-cost wearable, there is no guarantee that it will work with their smartphone, as ecosystems have unique compatibilities due to ”proprietary conditions”, and replacing wearables becomes an expensive affair.
Our interview data further supports this observation, as 10/22 participants only used their PHI wearable because it was gifted, and 3/7 participants who stopped using their PHI wearable did so due to incompatible ecosystems. As technology advances, it is clear that a unified framework for data storage and processing of PHI is possible; however, the current state of affairs indicates that factors unrelated to feasibility (market share and profitability) hinder this progress.
Another aspect of fragmentation that affects user adoption is the lack of reliable systems supporting collective care. In India, care (especially for health and wellbeing) is enacted not individually or by self, but as a collective. This was observed from the participants’ motivation/purpose for using PHI, as noted in both the interviews and the survey. Even those with a focus on self-improvement did so with the need to care for their loved ones in mind. This further indicates that even if a unified ecosystem is developed, it will be lacking in support for care realities in India unless it encompasses all stakeholders involved.
”I usually share health reports with family members as and when we get health checkups done, however, they usually get scattered all over in phone storage. If they can be stored all together in one place, that would be a helpful feature. ” - Survey Respondent (F, 33)
4.1.2. Low digital health literacy
At the individual level, a few participants found their level of digital health literacy to be a barrier to sustained usage of PHI. Raw metrics and clinical jargon were frequently reported as obstacles to understanding. Users often receive numbers without clear next steps. While numbers were sufficient for simpler metrics, the difficulty in taking action using derived advanced metrics proved to be a hurdle. Sometimes, users were not even aware of the possible features due to a lack of understanding or because the features were less popular among those without experience in self-tracking. Low digital health literacy, therefore, blocks interpretation and action. For instance, participant P10 admitted that he did not understand whether a certain value of heart rate was beneficial for him or not; however, he trusted the recommendations from doctor visits, which suggested what was good and what was observed during the visit.
Participants asked for plain-language explanations and threshold-based “what to do” guidance tied to their readings, as P9 noted through her personal experience. These needs emerged repeatedly throughout our study and motivate the development of literacy-appropriate microcopy and tooltips that are actionable not only to the user but also to the stakeholders involved.
4.1.3. Trust and sensor-confidence (cross-device checks)
As the other side of the coin of digital health literacy, lack of confidence in sensors emerged as a critical barrier. Participants routinely cross-checked device readings and sought confirmation from clinicians or trusted experts before taking action. Conditioned trust was common, treating wearable outputs as indicative but not definitive. Some attributed this lack of inherent trust also to a lack of familiarity with using these new-age devices, as they had grown up with different devices (such as pressure cuffs, as observed by P5).
Doubts about the accuracy of derived metrics (e.g., advanced sleep stages or spO2 estimates) led to hesitancy in taking preventive action, and participants suggested that the lack of confidence indicators or provenance metadata contributed to their reluctance to act on device signals. This is evident from P4’s notion of ”not blindly believing” in PHI systems when they indicate that something has gone wrong.
While this observation suggests that improving digital health literacy may overcome this barrier, the current state-of-the-art PHI systems used in Indian contexts still rely on black-box models for explaining readings and do not provide provenance for the insights delivered. The lack of infrastructure to support this logging, combined with the limited digital health education programs in India, further exacerbates this barrier.
4.1.4. Device/form-factor friction
While fragmentation, low health literacy, and a lack of trust were prevalent, physical discomfort (skin irritation), battery demands, and simple inconvenience led other participants to abandon continuous wearing. For example, while smartwatches are the most common form of PHI wearables, P22’s discomfort with wearables stems from her allergic reaction to straps being worn for long durations. While other form factors (in the form of rings and pendants) have started to emerge in the market, the lack of economic feasibility limits the utilization of such PHI systems.
Form-factor constraints, therefore, interrupt the continuity of sensing and undermine trend detection. Several participants expressed a preference for alternatives (phone-based sensing, rings, or non-wrist wearables) and suggested design changes (battery-life indicators, low-sampling “always-on” modes) to reduce attrition or to continue wearing other watches (as in the case of P19).
4.1.5. Preference for embodied and traditional cues (reactive norms)
While the previous barriers had external influences, this one is particularly dependent on the users’ state of mind. A strong normative tendency to rely on felt symptoms and clinic visits reduces engagement with preventive PHI. For example, P20’s notion of taking medication or precaution if he felt something was wrong indicates his strong trust in self-perception about his wellbeing.
”If there is something wrong with my body, maybe, I know or maybe my body will let me know if I’m feeling something bad. So in that case I will take medication or precaution or something, whatever that is needed at that time.” - P20 (M,29)
For others, embodied cues and clinician validation remain the primary triggers for action; continuous sensor outputs are viewed as supplemental rather than decisive. For example, P15 emphasized the need to integrate cultural cues into understanding the state of the body, rather than relying solely on continuous monitoring of vital signs. This cultural and behavioral preference helps explain why PHI often remains reactive in practice.
”If you think about it, our lifestyles has been completely different from our parents. These (PHI) trackers only work for us. If they are able to integrate regular daily activities (like cleaning the house, washing clothes, gardening) as part of measuring other metrics for health markers, I believe it can improve the adoption and create a more proactive health management mindset.” - P15 (F,25)
As Indians slowly shift towards a proactive mindset, the tools to support them are lacking, as most are designed with the mindset of supporting WEIRD contexts. The lack of equitable access and poor alignment with cultural and societal norms remain critical barriers to the widespread adoption of PHI for proactive collective care.
4.2. Enablers of Proactivity: Behaviours & Triggers
While barriers focused on why PHI was not being used effectively for proactive care, participants also highlighted several behaviors and enabling mechanisms that encourage earlier action and sustained monitoring.
4.2.1. Stakeholder specific sharing
In our sample, the family served as the primary channel for the introduction, mediation, and prompting of PHI. Younger family members often set up devices for older adults. For example, P8 wanted to keep an eye on his parents, who lived far away from him, and decided to monitor them through Fitbits.
Even during the co-design workshops and stakeholder interviews, family requests or approvals were repeatedly cited as immediate triggers for checking metrics or seeking care. Role-based sharing (care circles with roles such as monitor, nudge, and emergency) was widely seen as useful, provided that sharing is granular and consented. These findings underscore the importance of family workflows as both adoption and proactive levers. For example, a couple (CP7 and CP8) noted that as long as they did not have to manually add every granular detail, a dedicated platform for family-based sharing was beneficial.
Beyond sharing with family (and friends in some cases), participants expressed a strong desire for concise, clinician-ready summaries that could be shown at visits. The prospect of clinician engagement increased the perceived credibility of PHI and would motivate sustained monitoring. Several respondents said they would act earlier if a clinician requested or validated specific metrics. In the co-design workshop, another couple (CP6 and CP5) discussed their experience from the UK, using examples from a management-centric application with their GP, and mentioned that having control over sharing summaries with stakeholders, similarly based on their needs, would improve engagement.
@@ 4.2.1. Stakeholder specific sharing
While literature and current practices suggest that such sharing mechanisms exist specifically for each hospital to share information, such as clinical reports or lab test results, the ability to share PHI summaries with doctors or other critical stakeholders remains lacking.
4.2.2. Actionable explanations and engagement
Beyond the aspect of sharing information with stakeholders, plain-language recommendations, threshold-based guidance, and short “next-step” prompts (e.g., “re-measure/share with clinician/rest & re-check in 30 min”) were widely requested. Actionable microcopy converts numbers into behavior, reducing interpretation burden and increasing the likelihood of timely response. For example, participant P7’s experience of how notifications from his watch motivated him to walk more indicates that actionable insights are useful motivators. Participants welcomed light-touch social features as mechanisms to sustain routine monitoring. This included family challenges, badges, and progress nudges, as suggested by CP1, as well as viewing statistics on a routine basis in a graphical format, as suggested by CP3.
@@ 4.2.2. Actionable explanations and engagement
Participants noted that having a healthy competition among family members would push each other to strive for a healthier lifestyle, indicating that such features serve both as positive engagement and action-triggering functions. They create repeated opportunities for attention and for family-led prompting that can convert sensor signals to action towards a proactive, healthy life.
4.2.3. Proactive Monitoring vs Reactive Verifying
With the broader analysis, participants’ engagements with PHI fell broadly along a proactive–reactive axis. A subset of respondents described routine, preventive behaviors such as regular step-counting (as in P16), sleep checks, or goal-driven monitoring, which we label as proactive monitoring.
@@ 4.2.3. Proactive Monitoring vs Reactive Verifying
Others reported episodic use of PHI, typically in response to symptoms or clinical prompts, which we label reactive use. These two modes coexist. Many participants maintained short-term, goal-oriented monitoring while reserving clinician contact for symptomatic events. This pattern explains why PHI often fails to translate into early clinical action despite widespread adoption of sensing devices in the sample.
Participants described several triggers that convert passive sensing into action. Broadly, from the co-design sessions, social prompts (family requests and shared alerts), clinician-initiated requests, and interface-level cues (approval flows, explicit “ask to share” prompts) with granular user controls (as suggested by CP1) were repeatedly reported as immediate motivators for re-checking or seeking care. These observations motivated our operational focus. A PHI system must simultaneously enable agency (who controls the data), manage elicitation (confidence indicators and escalation paths) to shift use toward proactivity, and provide engagement (actionable explanations, social workflows).
4.3. Prototyping & Evaluation
Based on all the data from the previous components of the study, we designed and evaluated a medium-fidelity prototype focused on sharing preferences, family controls, and summary panels. Participants evaluated the prototype’s modular sharing UI, approval flows, and visualization options. The primary screen’s UI was designed to provide contextualized, actionable advice along with common physiological values. The sharing preferences were designed to be granular and customizable per stakeholder involved. Utilizing the critical elements discussed during the interviews and co-design workshops, which focused on the balance between control and ease of access, the buttons were designed to share health features among care circle members. Selected screens are shown in Figure 3. The takeaways from the prototype evaluation are listed in Table 4.
| Likes | Respondents appreciated the simplicity and modularity of the “Data & Sharing” center, the ability to craft templates (e.g., “weekly summary to family; alerts to clinician only”), and the “ask to share” approval flow. Graphical daily/weekly/monthly summaries were widely preferred for trend recognition. |
| Dislikes/change requests | Participants sought clearer, plain-language explanations (what does this sleep score mean?), clinician export templates, and EHR export options. They also requested confidence indicators explaining why a reading should (or should not) be trusted. |
| Perceived Benefits | Several participants stated that the prototype’s granular sharing and prompt features would likely cause them to act earlier (e.g., prompting a parent to visit a clinician after a flagged trend). Thus, while not a field evaluation, the prototype functioned as a proof-of-concept, validating several design implications (agency controls, confidence badges, clinician summaries). |
5. Discussion and Design Implications